solve the equation x^3+67x-126=0 and pls give a generalised method of solving cubics, 


 thank u

Grade:11

3 Answers

Jaya IITK
12 years ago

theres no method of solving cubic eqn u have to put values like 1,2,-1,-2 etc and find out the most appropriate one

 

BEST OF LUCK!!!!!!!!!!!!!!!!

 

Pls approve my answer by clicking on yes below!!!!!!!!!

SPARSH JAUHARI
12 years ago

i think i doesn't have any solution.Generally in cubic we try to get factors that make eqn=0
Then we solve by factorising it.But sometimes we  should put the values of x as 1,-1,2,-2,3,-3.Mostly you will get an  idea which value to put.
Like here i put x=2
we get 8+134-126 but that's not = to  0 .  
putting 1or -1 will not make a total to cancel 126. similarly if you'll put values 3,-3 or greater than that,you'll get total much above 126,so you'll not be able to cancel.
 Now take the equation: 

 

x^3-67x+126=0 (different than the question one)
here if you put x=2  you'll get 8-134+126=126-126=0


This is the method i want to convey it is helpful at times when you are not able to get by factorisation method.
ALL  THE BEST 


 
                                     

samlesh choudhury
12 years ago

its easy just pu thte factors of the constant then u wil get the root